OpenStudy (anonymous):

With respect to the origin O, the points A and B have position vectors given by −−→OA = i + 2j + 2k and −−→OB = 3i + 4j. The point P lies on the line AB and OP is perpendicular to AB. (ii) Find the position vector of P.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

I like to use <a,b,c> to represent a vector rather than a i + b j + c k. (easier to type) A is at (1,2,2) B is at (3,4,0) the vector from A to B is B-A. I would find B-A

OpenStudy (anonymous):

then

OpenStudy (phi):

write down the equation of the line AB P (for point on the line) = <1,2,2> + t (B-A) where B-A is ?

OpenStudy (anonymous):

2,2,-2

OpenStudy (phi):

ok, and because we only care about the direction we can write it as 2<1,1,-1> and ignore the 2 the equation of the line AB is P= <1,2,2> + t<1,1,-1> where t is a scalar

OpenStudy (anonymous):

and?

OpenStudy (phi):

I was thinking OP perpendicular to the line means means P is perpendicular to the direction vector B-A i.e. to <1,1,-1> and when we dot product P with <1,1,-1> we get 0

OpenStudy (anonymous):

and how do we do that

OpenStudy (phi):

let P= <x,y,z> then because P is orthogonal to <1,1,-1> we know <x,y,z> dot <1,1,-1> =0 and x + y - z =0

OpenStudy (phi):

using the equation of the line P= <1,2,2> + t<1,1,-1> <x,y,z> = <1+t, 2+t, 2-t> we see x= 1+t y= 2+t z= 2-t put those into x + y - z =0

OpenStudy (anonymous):

i dont think it is good ?

OpenStudy (phi):

what do you mean ?

OpenStudy (anonymous):

op = 2/3i + 5/3 j + 7/3 k

OpenStudy (anonymous):

this is the answer u should be getting

OpenStudy (phi):

yes, that is what we will get

OpenStudy (anonymous):

how ? i am not getting it

OpenStudy (phi):

do you understand this part: let P= <x,y,z> then because P is orthogonal to <1,1,-1> we know <x,y,z> dot <1,1,-1> =0 and x + y - z =0

OpenStudy (anonymous):

yes

OpenStudy (phi):

also, P is on the line so it must be that P= <1,2,2> + t<1,1,-1> <x,y,z> = <1+t, 2+t, 2-t>

OpenStudy (anonymous):

yes good

OpenStudy (phi):

and component by component, <x,y,z> = <1+t, 2+t, 2-t> means x= 1+t y= 2+t z= 2-t

OpenStudy (anonymous):

yes

OpenStudy (phi):

now put all the variables in terms of "t" in x + y - z =0 using x= 1+t y= 2+t z= 2-t and solve for t

OpenStudy (anonymous):

t is ?

OpenStudy (phi):

the equation of a line in n-dim is typically written as \[ P= P_0 + t \vec{V} \] where V is the direction we move in starting at some point P0 the "t" is a scalar that we multiply the direction vector by to show how far we move we are using the info in this problem to find the "t" that gets us to a point that solves the problem

OpenStudy (anonymous):

i mean how much t is ?

OpenStudy (anonymous):

what answer uve got for t

OpenStudy (phi):

x + y - z =0 using x= 1+t ,y= 2+t ,z= 2-t replace x with 1+t, etc (1+t) + (2+t) - (2-t) = 0

OpenStudy (anonymous):

yes i know this

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

-1/3

OpenStudy (phi):

now use P= <x,y,z> = <1,2,2> + t<1,1,-1> = <1+t, 2+t, 2-t> to find the answer

OpenStudy (anonymous):

is t -1/3 ?

OpenStudy (phi):

and then write it using i, j and k yes , t= -1/3

OpenStudy (phi):

if you do the problem with t= -1/3 you will get the answer you posted earlier.